# The moduli spaces of left-invariant statistical structures on Lie groups

### v1

@@ Abstract.
The first author is supported by JST SPRING, Grant Number JPMJSP2132. The second author is supported by JST SPRING, Grant Number JPMJSP2119. The third author is supported by JSPS KAKENHI Grant Number JP24K06714. The fourth author is supported by JSPS KAKENHI Grant Numbers 23K22395, 24K21193. This work was partly supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 and Osaka Metropolitan University Strategic Research Promotion Project (Development of International Research Hubs). This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@ 1. Introduction
We will see that two left-invariant statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ belongs to an $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$-orbit if and only if there exists a scaling statistical isomorphism compatible with the group structure (see Definition 4.2.1) between statistical Lie groups $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$. In particular, the action of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ preserves the properties of statistical structures, such as dually flatness and conjugate symmetry. Consequently, we define two subspaces of $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$, denoted by $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, as the moduli spaces of left-invariant conjugate symmetric statistical structures and left-invariant dually flat structures, respectively. In this paper, motivated by Problem A, we address the following problem:
@@
(1.1) $\mathbb{R}^{n},\quad G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2),\quad H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
@@
For these Lie groups $G$, let us denote by $S^{3}(\mathfrak{g}^{\ast})$ the space of all symmetric $(0,3)$-tensors on the Lie algebra $\mathfrak{g}$ of $G$. Then the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ can be written as an orbit space of $S^{3}(\mathfrak{g}^{\ast})$ under a certain Lie group (see Section 4.3). A set of representatives for the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$, resp. $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, refers to a subset of $S^{3}(\mathfrak{g}^{\ast})$ whose projection onto these moduli spaces is surjective. If this projection is bijective, we call it a complete set of representatives. The notations $V^{+}$, $v_{0}$, and $w_{0}$ appearing in Table 1 will be explained in Theorems 1.0.1, 1.0.2, and 1.0.3, respectively.
@@
Corollary 1.0.4 (see also Theorem 6.3.1).
@@ 2.2. Statistical manifolds
In this paper, we say that a triple $(M,g,\nabla)$ is called a statistical manifold if $g$ is a Riemannian metric and $\nabla$ is an affine connection on $M$ satisfying all conditions in Proposition 2.1.1. For a statistical manifold $(M,g,\nabla)$, the pair $(g,\nabla)$ is called a statistical structure on $M$. In particular, $\nabla$ is called a statistical connection on the Riemannian manifold $(M,g)$.
For each Riemannian metric $g$ on $M$, the pair $(g,\nabla^{g})$ is a statistical structure on $M$ by the definition of the Levi-Civita connection, and $(g,\nabla^{g})$ is called the trivial statistical structure (see [8]). Thus, statistical manifolds can be considered as a generalization of Riemannian manifolds.
@@
The following proposition will be applied in Theorem 2.3.4.
Proposition 2.2.2.
@@
By Proposition 2.2.3, a pair consisting of the Riemannian metric $g$ and a cubic form $C\in S^{3}(T^{\ast}M)$, or a pair consisting of $g$ and a difference tensor $K\in\mathcal{K}(M,g)$, can also be called a statistical structure on $M$. Thus, we will also call $(g,C)$ and $(g,K)$ statistical structures on $M$ as well. We see that the statistical structure $(g,0)$ on $M$ corresponds to the statistical structure $(g,\nabla^{g})$. Here 0 is a origin of $\mathcal{K}(M,g)$ or $S^{3}(T^{\ast}M)$ as the liner spaces.
@@
Throughout this paper, for each $K\in\mathcal{K}(M,g)$ and each vector field $X$ on $M$, we define the $(1,1)$-tensor field $K_{X}$ on $M$ by putting $K_{X}(Y):=K(X,Y)$ for each vector field $Y$. For a pair $X,Y$ of vector fields on $M$, the Lie bracket $K_{X},K_{Y}]$ of $K_{X}$ and $K_{Y}$ is defined as $K_{X},K_{Y}](Z)=K_{X}(K_{Y}(Z))-K_{Y}(K_{X}(Z))$. The following proposition, proved by Opozda [30, Corollary 3.8], will be applied in Section 5.2:
Proposition 2.2.4.
@@ A $(1,3)$-tensor field $\nabla^{g}K$ is totally symmetric, that is,
Note that, by Proposition 2.3.1, the following set is a linear subspace of $S^{3}(T^{\ast}M)$:
(2.3) $\left\{C\in S^{3}(T^{\ast}M)\ |\ (g,C)\text{ is conjugate symmetric}\right\}.$
@@ 2.4. Some examples of statistical manifolds
Example 2.4.1 ($n$-variable normal distribution family).
@@
[…]
@@
Moreover, a statistical connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}^{n},g^{F})$ correspondence to $(g^{F},C^{A(\alpha)})$ in the sense of Proposition 2.2.3 is called the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$.
@@
Example 2.4.3 (Takano Gaussian space).
@@
$\mathbf{p}_{T}=\mathbf{p}_{N}\circ\iota:\mathcal{N}^{n}_{T}\to\mathcal{P}(\mathbb{R}^{n}),\ \ (\Sigma,\mu)\mapsto N(x\mid\mathrm{diag}(\Sigma,\dots,\Sigma),\mu)dx.$
@@
$\iota:\mathcal{N}^{n}_{T}\hookrightarrow\mathcal{N}^{n},\ \ (\Sigma,\mu)\mapsto(\mathrm{diag}(\Sigma,\dots,\Sigma),\mu).$
Then the triple $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$ is a statistical model in the sense of [4]. On $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$, the Fisher metric $g^{F}$, the Amari–Chentsov $\alpha$-tensor field $C^{A(\alpha)}$, and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ are defined in the similar way as in Example 2.4.1, respectively. In [7], the Riemannian manifold $(\mathcal{N}_{T}^{n},g^{F})$ is called the Takano Gaussian space, and we use this terminology in this paper as well.
Remark 2.4.4.
@@ 3.1. Homogeneous statistical manifolds
Example 3.1.1.
The statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.1) is an $\mathrm{Aff}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the isotropy subgroup of the $\mathrm{Aff}(n,\mathbb{R})$-action on $\mathcal{N}^{n}$ at the point $(I_{n},0)$ is $O(n)\ltimes\{0\}$. Here, $I_{n}$ denotes the identity matrix of size $n$.
Example 3.1.2.
Let $\mathrm{Aff}^{d+}(n,\mathbb{R})\coloneqq\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$. The statistical manifold $(\mathcal{N}_{T}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.3) is an $\mathrm{Aff}^{d+}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the $\mathrm{Aff}^{d+}(n,\mathbb{R})$-action on $\mathcal{N}_{T}^{n}$ is simply-transitive.
@@ 3.2. Left-invariant Riemannian metrics and left-invariant statistical structures on Lie groups
  • $\mathcal{LA}_{\mathrm{Stat}}(G,g):=\{\nabla\in\mathcal{A}_{\mathrm{stat}}(G,g)\mid\text{$\nabla$ is left-invariant}\}\subset\mathcal{A}_{\mathrm{stat}}(G,g)$ and
  • $\mathcal{LK}(G,g):=\{K\in\mathcal{K}(G,g)\mid\text{$K$ is left-invariant}\}\subset\mathcal{K}(G,g)$.
@@
For the sake of computations in sections later, we summarize some correspondences in the form of the following proposition, which is a reformulation of Proposition 2.2.3.
Proposition 3.2.1.
@@
$\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ $\coloneqq\{(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)\mid(g,\nabla)\text{ is conjugate symmetric}\},$
@@
  • $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is conjugate symmetric}\}$,
  • $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is dually flat}\}$.
@@ 3.3. Some remarks on computations
where $\nabla^{(g,C)}$ denotes the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1. By Proposition 2.3.2, we shall regard the left-invariant $(0,4)$-tensor field $R^{C}_{g}$ on $G$ as an element of $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$, where $\wedge^{2}\mathfrak{g}^{\ast}$ denotes the space of alternating $2$-tensors on $\mathfrak{g}^{\ast}$, and $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ the space of symmetric $2$-tensors on $(\wedge^{2}\mathfrak{g}^{\ast})$. Let us put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$ for each $i,j$. Then $\{\omega_{ij}\}_{i<j}$ forms a basis of the space $\wedge^{2}\mathfrak{g}^{\ast}$ of alternating $2$-tensors on $\mathfrak{g}^{\ast}$. We shall write $(i,j)\leq(k,l)$ if $i<k$ or “$i=k$ and $j\leq k$”. Then the family $\{\omega_{ij}\odot\omega_{kl}\}_{i<j,k<l,(i,j)\leq(k,l)}$ forms a basis of linear space $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$.
@@
Throughout this paper, for each $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$, we denote the $(0,4)$-tensor $R_{g}^{C}-R_{g}^{0}\in(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ by $K^{(g,C)},K^{(g,C)}]$, or simply $[K,K]$ when the context is clear. Then by Proposition 2.3.2,
@@
$[K^{(g,C)},K^{(g,C)}]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij}),$

### v3

@@ Abstract.
The first author was supported by JST SPRING, Grant Number JPMJSP2132. The second author was supported by JST SPRING, Grant Number JPMJSP2119. The third author was supported by JSPS KAKENHI, Grant Number JP24K06714. The fourth author was supported by JSPS KAKENHI Grant Numbers 23K22395, 24K21193. This work was partially supported by the MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 and the Osaka Metropolitan University Strategic Research Promotion Project (Development of International Research Hubs). This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center at Kyoto University.
@@ 1. Introduction
We will see that two left-invariant statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ belongs to an $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$-orbit if and only if there exists a scaling statistical isomorphism compatible with the group structure (see Definition 4.3) between statistical Lie groups $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$. In particular, the action of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ preserves the properties of statistical structures, such as dually flatness and conjugate symmetry. Consequently, we define two subspaces of $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$, denoted by $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, as the moduli spaces of left-invariant conjugate symmetric statistical structures and left-invariant dually flat structures, respectively. In this paper, motivated by Problem A, we address the following problem:
@@
(1.1) $\mathbb{R}^{n},\;G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2),\;H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
@@
For these Lie groups $G$, let us denote by $S^{3}(\mathfrak{g}^{\ast})$ the space of all symmetric $(0,3)$-tensors on the Lie algebra $\mathfrak{g}$ of $G$. Then the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ can be written as an orbit space of $S^{3}(\mathfrak{g}^{\ast})$ under a certain Lie group (see Section 4.3). A set of representatives for the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$, resp. $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, refers to a subset of $S^{3}(\mathfrak{g}^{\ast})$ whose projection onto these moduli spaces is surjective. If this projection is bijective, we call it a complete set of representatives. The notations $V^{+}$, $v_{0}$, and $w_{0}$ appearing in Table 1 will be explained in Theorems 1.1, 1.2, and 1.3, respectively.
@@
Corollary 1.4 (see also Theorem 6.8).
@@ 2.2. Statistical manifolds
In this paper, we say that a triple $(M,g,\nabla)$ is called a statistical manifold if $g$ is a Riemannian metric and $\nabla$ is an affine connection on $M$ satisfying all conditions in Proposition 2.1. For a statistical manifold $(M,g,\nabla)$, the pair $(g,\nabla)$ is called a statistical structure on $M$. In particular, $\nabla$ is called a statistical connection on the Riemannian manifold $(M,g)$.
For each Riemannian metric $g$ on $M$, the pair $(g,\nabla^{g})$ is a statistical structure on $M$ by the definition of the Levi-Civita connection, and $(g,\nabla^{g})$ is called the trivial statistical structure (cf. [8]). Thus, statistical manifolds can be considered as a generalization of Riemannian manifolds.
@@
The following proposition will be applied in Theorem 2.9.
Proposition 2.3.
@@
By Proposition 2.4, a pair consisting of the Riemannian metric $g$ and a cubic form $C\in S^{3}(T^{\ast}M)$, or a pair consisting of $g$ and a difference tensor $K\in\mathcal{K}(M,g)$, can also be called a statistical structure on $M$. Thus, we will also call $(g,C)$ and $(g,K)$ statistical structures on $M$ as well. We see that the statistical structure $(g,0)$ on $M$ corresponds to the statistical structure $(g,\nabla^{g})$. Here 0 is a origin of $\mathcal{K}(M,g)$ or $S^{3}(T^{\ast}M)$ as the liner spaces.
@@
Throughout this paper, for each $K\in\mathcal{K}(M,g)$ and each vector field $X$ on $M$, we define the $(1,1)$-tensor field $K_{X}$ on $M$ by putting $K_{X}(Y):=K(X,Y)$ for each vector field $Y$. For a pair $X,Y$ of vector fields on $M$, the Lie bracket $K_{X},K_{Y}]$ of $K_{X}$ and $K_{Y}$ is defined as $K_{X},K_{Y}](Z)=K_{X}(K_{Y}(Z))-K_{Y}(K_{X}(Z))$. The following proposition, proved by Opozda [29, Corollary 3.8], will be applied in Section 5.2:
Proposition 2.5.
@@ A $(1,3)$-tensor field $\nabla^{g}K$ is totally symmetric, that is,
Note that, by Proposition 2.6, the following set is a linear subspace of $S^{3}(T^{\ast}M)$:
(2.3) $\left\{C\in S^{3}(T^{\ast}M)~\middle|~(g,C)\text{ is conjugate symmetric}\right\}.$
@@ 2.4. Some examples of statistical manifolds
Example 2.10 ($n$-variable normal distribution family).
@@
$\mathbf{p}_{N}:\mathcal{N}^{n}\to\mathcal{P}(\mathbb{R}^{n}),~(\Sigma,\mu)\mapsto N(x\mid\Sigma,\mu)dx:=\frac{1}{\sqrt{(2\pi)^{n}\det(\Sigma)}}\exp\bigg(-\frac{1}{2}(x-\mu)^{\mathsf{T}}\Sigma^{-1}(x-\mu)\bigg)dx$
@@
Moreover, a statistical connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}^{n},g^{F})$ correspondence to $(g^{F},C^{A(\alpha)})$ in the sense of Proposition 2.4 is called the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$.
@@
Example 2.12 (Takano Gaussian space).
@@
$\mathbf{p}_{T}=\mathbf{p}_{N}\circ\iota:\mathcal{N}^{n}_{T}\to\mathcal{P}(\mathbb{R}^{n}),~(\Sigma,\mu)\mapsto N(x\mid\mathrm{diag}(\Sigma,\dots,\Sigma),\mu)dx.$
@@
$\iota:\mathcal{N}^{n}_{T}\hookrightarrow\mathcal{N}^{n},~(\Sigma,\mu)\mapsto(\mathrm{diag}(\Sigma,\dots,\Sigma),\mu).$
Then the triple $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$ is a statistical model in the sense of [4]. On $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$, the Fisher metric $g^{F}$, the Amari–Chentsov $\alpha$-tensor field $C^{A(\alpha)}$, and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ are defined in the similar way as in Example 2.10, respectively. In [7], the Riemannian manifold $(\mathcal{N}_{T}^{n},g^{F})$ is called the Takano Gaussian space, and we use this terminology in this paper as well.
Remark 2.13.
@@ 3.1. Homogeneous statistical manifolds
Example 3.1.
The statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.10) is an $\mathrm{Aff}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the isotropy subgroup of the $\mathrm{Aff}(n,\mathbb{R})$-action on $\mathcal{N}^{n}$ at the point $(I_{n},0)$ is $O(n)\ltimes\{0\}$. Here, $I_{n}$ denotes the identity matrix of size $n$.
Example 3.2.
Let $\mathrm{Aff}^{d+}(n,\mathbb{R})\coloneqq\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$. The statistical manifold $(\mathcal{N}_{T}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.12) is an $\mathrm{Aff}^{d+}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the $\mathrm{Aff}^{d+}(n,\mathbb{R})$-action on $\mathcal{N}_{T}^{n}$ is simply-transitive.
@@ 3.2. Left-invariant Riemannian metrics and left-invariant statistical structures on Lie groups
  • $\mathcal{LA}_{\mathrm{Stat}}(G,g):=\left\{\nabla\in\mathcal{A}_{\mathrm{stat}}(G,g)~\middle|~\text{$\nabla$ is left-invariant}\right\}\subset\mathcal{A}_{\mathrm{stat}}(G,g)$ and
  • $\mathcal{LK}(G,g):=\left\{K\in\mathcal{K}(G,g)~\middle|~\text{$K$ is left-invariant}\right\}\subset\mathcal{K}(G,g)$.
@@
For the sake of computations in sections later, we summarize some correspondences in the form of the following proposition, which is a reformulation of Proposition 2.4.
Proposition 3.3.
@@
$\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ $\coloneqq\left\{(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)~\middle|~(g,\nabla)\text{ is conjugate symmetric}\right\},$
@@
  • $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g):=\left\{C\in S^{3}(\mathfrak{g}^{\ast})~\middle|~(g,C)\text{ is conjugate symmetric}\right\}$,
  • $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g):=\left\{C\in S^{3}(\mathfrak{g}^{\ast})~\middle|~(g,C)\text{ is dually flat}\right\}$.
@@ 3.3. Some remarks on computations
where $\nabla^{(g,C)}$ denotes the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.3. By Proposition 2.7, we shall regard the left-invariant $(0,4)$-tensor field $R^{C}_{g}$ on $G$ as an element of $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$, where $\wedge^{2}\mathfrak{g}^{\ast}$ denotes the space of alternating $2$-tensors on $\mathfrak{g}^{\ast}$, and $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ the space of symmetric $2$-tensors on $(\wedge^{2}\mathfrak{g}^{\ast})$. Let us put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$ for each $i,j$. Then $\{\omega_{ij}\}_{i<j}$ forms a basis of the space $\wedge^{2}\mathfrak{g}^{\ast}$ of alternating $2$-tensors on $\mathfrak{g}^{\ast}$. We shall write $(i,j)\leq(k,l)$ if $i<k$ or “$i=k$ and $j\leq k$”. Then the family $\{\omega_{ij}\odot\omega_{kl}\}_{i<j,k<l,(i,j)\leq(k,l)}$ forms a basis of linear space $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$.
@@
Throughout this paper, for each $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$, we denote the $(0,4)$-tensor $R_{g}^{C}-R_{g}^{0}\in(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ by $K^{(g,C)},K^{(g,C)}]$, or simply $[K,K]$ when the context is clear. Then by Proposition 2.7,
@@
$K^{(g,C)},K^{(g,C)}]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij}),$